OpenStudy (anonymous):

O2 can be produced by the reation 2 KCI3 --- 2KCI + 3O2. In one experiment .312 g of O2 were formed. How much KCIO3 was decomposed?

OpenStudy (anonymous):

use equv law \[m(KClO _{3})/M _{e}(KClO _{3})=m(O _{2})/M _{e}(O _{2})\] the Me(O2)= 8g/mol , Me(KClO3)=122,5 g/mol so you will get m(KClO3)=4,7775g